the first of two numbers is 3 times the second.thier sum is 88. find the numbers
Answer by oberobic(2304)

22 + 66 = 88
.
But how do you get that?
.
The sum of two numbers is 88
x+y = 88
.
The first is 3 times the second
x = 3*y
.
Subsitute
.
3y + y = 88
4y = 88
y = 22
.
x + y = 88
x = 88-y
x = 88-22
x = 66
.
Done.